# Title keyword 5 for citation Title for citation Title for citation

### v1

@@ 1.1 Requirements and Challenges in the Post-5G Era
1.1.1 Larger Bandwidth and Better Utilization of Spectrum Resources
The rapid development of information technology and networks has led to a new round of business innovation and upgrading of service demand. The networks face the dual challenges of explosive growth in data traffic and the demand for massive devices to connect everything. According to Cisco statistics [2], global mobile data traffic grew more than tenfold from 2014 to 2019, and the number of connected devices worldwide had reached approximately 29.3 billion by the end of 2023. The communication networks and infrastructure must continuously improve their bandwidth capacity to satisfy users’ demand for high-speed and large-capacity data transmission. As an essential mobile communications resource, spectrum remains the most critical element for enhancing network capacity and ensuring that networks continue to provide high-quality communications services [3]. On the one hand, wireless networks are in dire need of tapping new spectrum resources as the sub-6GHz band is very crowded. On the other hand, existing wireless communication systems still adopt a static spectrum management model, where authorized users have exclusive access. This model struggles to flexibly adapt to the significant spatiotemporal variations in traffic and the mobility characteristics of users, resulting in low spectrum resource utilization.
1.1.2 Lower-Carbon Wireless Coverage
With the widespread adoption and continuous development of 5G networks, issues related to energy consumption and costs are becoming increasingly prominent. According to the 2020 White Paper on Energy-Saving Technologies for 5G base stations released by China Mobile [4], the typical power consumption of a 5G base station (BS) is about 3 to 5 kilowatts, which is three to four times that of a comparable 4G BS. Additionally, due to the use of higher frequency bands in 5G technology, there is a further reduction in the coverage area per BS. To achieve the same coverage targets, the number of 5G BSs required would be three to four times that of 4G, significantly increasing the cost of network deployment for 5G BSs. Moreover, the current operational management systems of actual networks lack flexibility and intelligence. The operational status of BSs cannot be dynamically adjusted in precise accordance with real-time changes in regional user traffic. During periods and in areas with low traffic demand, the BSs may maintain a high workload, leading to energy waste. In recent years, computing power, as a new form of productivity, has been driving the rapid growth of the digital economy. Data centers, serving as the physical infrastructure for computing power, inevitably exacerbate high energy consumption issues due to their large-scale construction and operation.
1.1.3 More Efficient and Cost-effective Network Operation, Maintenance, and Management
@@
Compared to traditional core network equipment’s typical millisecond-level fault detection efficiency, network resource virtualization introduces new potential fault points. Moreover, it is difficult for operators to delineate the responsibility boundaries among suppliers when faults occur. Using software mechanisms for fault detection, the virtual network function at the edge site typically experiences longer response times, which complicates timely network maintenance and management. Furthermore, the continuous emergence of new applications and services demands higher bandwidth, lower latency, more excellent reliability, and enhanced security from the network. Traditional network operation and maintenance management methods will find it increasingly difficult to adapt to the future’s more complex and dynamic network environments.
@@ 2.1.2 Deep Learning
The principle of deep learning technology is to emulate the human brain [36], constructing artificial neural networks to learn knowledge from data. The neural layers can automatically learn and extract complex feature patterns from large-scale datasets, using gradient descent to update parameter errors. Table 4 summarizes some popular deep learning algorithms, and detailed descriptions are given as follows. Convolutional neural networks (CCNs) are widely used for image and video understanding [37]. Inspired by the functioning principles of the biological visual system, CNNs efficiently extract local features from image data via convolution operations and reduce computational complexity through pooling layers. Recurrent Neural Networks (RNN) represent a class of neural networks designed explicitly for sequential data within deep learning. Leveraging a recursive network structure[38], RNNs excel at capturing latent dependencies in sequential data. Generative adversarial networks (GANs) are a class of neural networks widely used for unsupervised learning. They comprise two neural networks, a discriminator and a generator, and use adversarial training to produce artificial data identical to actual data [39]. Deep Reinforcement Learning (DRL) is a convergence of deep learning and reinforcement learning[35], employing deep neural networks to approximate value functions or policy functions in reinforcement learning problems. This approach enables agents to autonomously learn and optimize decision-making processes to maximize long-term rewards in complex and dynamic environments[40, 41].
@@
CNN LeNet-5 Early convolutional neural networks
@@ 3.2.2 AI-Based BS Energy Conservation
AI strategy can quickly analyze large-scale energy consumption data and discover hidden rules and correlations [85]. Therefore, network traffic can be accurately predicted according to dynamic data analysis results to reasonably plan parameters such as BS power and hibernation state and achieve accurate energy consumption control and resource optimization, as shown in Figure 2. The AI method can also systematically consider many factors, such as network coverage, capacity, and signal quality, to dynamically adjust the energy-saving strategy in the network and realize the global optimization decision. In addition, primary data such as the configuration and performance statistics of the live network can also be used to optimize the 5G device-level energy-saving function parameters based on feature clustering and neural network algorithms. Under the premise of ensuring service quality, the problem of 4G/5G overlapping coverage can be fully considered for the cell’s cooperative shutdown to maximize the effect of the device-level energy conservation function [86].
@@ 3.3.1 AI-Based Channel Estimation
Considering the sparse mmWave channel matrix as a natural image, the authors of [106] propose a practical and accurate channel estimation framework based on the fast and flexible denoising convolutional neural network. Simulation results validate that the training speed of a quick and flexible denoising convolutional neural network is more rapid than that of other channel estimators without sacrificing normalized mean square error performance. [107] considered the time-frequency response of a fast-fading communication channel as a 2D image. This scheme considers the pilot values as low-resolution images and uses an SR network cascaded with a denoising IR network to estimate the channel. The results confirm that this pipeline can be used efficiently in channel estimation. To address channel distortion, the model developed in [102] is first trained offline using the data generated from simulation based on channel statistics and then used for recovering the online transmitted data directly. [102] is more robust than conventional methods when fewer training pilots are used, the cyclic prefix is omitted, and nonlinear clipping noise exists. The authors of [108] exploited a learned denoising-based approximate message passing network, which can learn channel structure and estimate channels from many training data. Simulation results show that this neural network significantly outperforms state-of-the-art compressed sensing-based algorithms even when the receiver is equipped with a small number of RF chains.
@@ 3.3.2 AI-Based OFDM Receiver
The wireless communication transmitter and receiver are divided into multiple processing mods in OFDMules. Each module is responsible for specific sub-tasks, such as Quadrature amplitude modulation, an IFFT, channel estimation, signal detection, etc. The advantage of this design is that each module can be analyzed and optimized individually, resulting in a very efficient and stable system.
@@
In [109], a model-driven deep learning (DL) approach was proposed that combines deep learning with expert knowledge to replace existing orthogonal frequency division multiplexing receivers in wireless communications. In [110], a machine learning (ML) assisted physical layer receiver technique was proposed to demodulate the OFDM signals, subject to very high Doppler effects and corresponding distortions in the received signal. To find a balance between full-size CP and non-existent CP, the authors of [111] investigate the redundancy problem and propose a minimum redundant OFDM receiver using deep learning (DL) tools. In [112], the receivers were designed based on deep neural networks (DNNS) in deep learning, consisting of a layer of DNNS and soft decisions. The problems of channel estimation error, delay, and decoding limits between users with traditional detection methods have been solved. It can recover the symbols of all users at one time and jointly perform channel estimation and signal detection. In [113], a novel generative supervised deep neural network is designed that can be trained using a reasonable number of pilots. After channel estimation, the neural network-based receiver (specifically, the autoencoder) jointly learns the pre-encoder and decoder for data symbol detection. In [114], an intelligent receiver for the OFDM communication system was designed based on the Densenet neural network structure, and the intelligent receiver is realized by optimizing the Densenet neural network structure. This method can recover information at the receiving end and avoid complicated pilot operations and signal error accumulation. In [115], a convolutional neural network (CNN) was used to reduce the bit error rate by using the mathematical calculation function of discrete Fourier transform (DFT) and the training of OFDM signal samples.
@@ 3.3.4 AI-Based CSI Feedback Algorithm
3.4 New Businesses Enabled
AI algorithms can dynamically allocate resources, such as bandwidth and power, to create and manage network slices in the wireless network. This can enable the creation of customized networks tailored to the needs of different applications and services. The booming development of AI has enabled the development of many new businesses, such as network slicing, mobility management, and wireless localization.
@@ 4.2.1 Distributed Intelligence and Federated Learning
The NET4AI architecture must support low-latency, high-reliability communication computing requirements and increasing privacy concerns. Distributed Intelligence and Federated Learning are the key technologies that can solve this problem. Traditional centralized architectures in wireless networks can no longer meet low-latency, high-reliability communication and computation demands. Additionally, conventional centralized architectures are inefficient in supporting the ubiquitous intelligence required by future wireless networks. This trend towards deploying intelligent decentralized elements has become apparent, with intelligence gradually shifting from centralized to distributed systems. Therefore, introducing distributed intelligent computing architectures is necessary to fully utilize the multi-dimensional data and computing resources held by user terminals and nodes, enabling intelligent connectivity.
@@
Figure 18: Federated Learning, Multi-Agent Reinforcement Learning and Split Learning
@@
Horizontal Federated Learning: Also known as sample-based federated learning, this can be applied in scenarios where the datasets of the various participants in federated learning have the same feature space but different sample spaces. Horizontal federated learning enhances privacy by training AI models on local devices and sharing only model parameters rather than raw data. Each node trains the model locally during training and periodically uploads model weights. The central node integrates these weights to form federated model weights, which are then sent back to the nodes for the next training cycle or local inference. Federated learning ensures that user data remains local, transmitting only weights/gradients to achieve global training, thereby enhancing privacy.
@@
Through horizontal federation learning, different participants can jointly train a more powerful AI model based on local data sets while protecting privacy and improving the accuracy and generalization of the model. Additionally, horizontal federated learning enables real-time learning on distributed edge devices, allowing for timely model parameter updates to constantly changing environments and data. This is crucial for AI applications that require quick responsiveness and strong adaptability. It also supports collaboration across devices with varying computational capabilities and data distributions, catering to the diverse needs of 6G network terminals. Personalized model training on local devices enhances user experience by offering more individualized services.
@@
The main application areas of horizontal federated learning include smart healthcare, personalized services, and intelligent manufacturing[181]. The authors of [182] provide the framework for an advanced blockchain-orchestrated machine learning system for privacy-preserving federated learning in medicine and a new utility in health. The authors of [183] combined blockchain and federated learning to accelerate AI model training in healthcare and improve model robustness. The authors of [184] introduced personalized federated learning (PFL) methods with theoretical performance derivations, modeling user personalization and data distribution. The authors of [185] categorized PFL techniques based on key challenges and strategies, highlighting critical ideas and future research directions in PFL architecture, realistic benchmarks, and trustworthy methods. The authors of [186] propose a new federated learning framework for cross-domain prediction for intelligent manufacturing, designed to quickly adapt to new products, processes, and applications with limited training data.
@@
Vertical Federated Learning: Vertical federated learning refers to a setup where datasets have different features but share the same samples/users. It is suitable for scenarios where there is a high overlap of users among participants but low or no overlap in feature space by extracting parts where participants have the same users but not the same features for joint training. In the training process, it is first necessary to sample align each participant’s data to obtain overlapping user data, then train the model locally and upload the parameters to the server [187]. The difference between vertical and horizontal federated learning is that horizontal federated learning emphasizes cooperation and sharing between different data sources. In contrast, vertical federated learning focuses on collaboration and sharing within varying levels of the same data source.
@@
Vertical federated learning can leverage the correlation between different levels of data sources to achieve data augmentation. By sharing model parameters, various data sources can complement and enrich each other’s data, improving AI models’ generalization ability and robustness. Furthermore, vertical federated learning can facilitate cross-level learning and knowledge transfer. Different data sources can share model parameters and learning experiences, achieving cross-level knowledge transfer and migration and enhancing the overall model’s effectiveness.
@@
The main application areas of vertical federated learning include recommendation systems, smart healthcare, and smart transportation. The authors of [188] propose a method based on clustering and latent factor models under the vertical federated recommendation system. Considering the diversity of a large number of different users in each participant, users are clustered to reduce the dimension of the matrix and improve the accuracy of user recommendations. The authors of [189] describe the idea of Intention-Hiding Vertical Federated Learning and illustrate a framework to achieve this privacy-preserving goal. The authors of [190] propose a Federated Deep Learning based on the Spatial-Temporal Long and Short-Term Networks algorithm to predict traffic flow by utilizing observed historical traffic data. An additive homomorphic encryption method based on Vertical Federated Learning trains local TFP models and shares short-term spatio-temporal hiding information.
@@
Split Learning: This technique separates machine learning model architecture between clients and servers, allowing joint model training without disclosing raw data. High-computation tasks are assigned to powerful central nodes, while shallower data processing layers are kept at data-holding terminals. Split learning offers better privacy than federated learning and reduces the processing load on resource-limited devices. Its applications focus on healthcare, where lightweight devices train intelligent models based on users’ private health data [197].

### v2

@@ 1.1 Requirements and Challenges in the Post-5G Era
1.1.1 More Efficient Resource Utilization
The rapid development of information technology and networks has led to a new round of business innovation and upgrading of service demand. The networks face the dual challenges of explosive growth in data traffic and the demand for massive devices to connect everything. According to Cisco statistics [2], global mobile data traffic grew more than tenfold from 2014 to 2019, and the number of connected devices worldwide had reached approximately 29.3 billion by the end of 2023. In the face of such an enormous traffic demand, it is of particular significance to explore how to utilize the limited network resources as fully as possible [3]. However, existing wireless communication systems still adopt a static spectrum management model, where authorized users have exclusive access. This model struggles to flexibly adapt to the significant spatiotemporal variations in traffic and the mobility characteristics of users, resulting in low spectrum resource utilization, as well as a huge power consumption of the base stations. Consequently, in the event that the network is capable of, by virtue of the intelligent prediction and dynamic adjustment capabilities furnished by AI, allocating network resources such as spectrum and power in a flexible manner in accordance with the real-time demands of various services, it will be feasible to curtail resource wastage or the scarcity of resources for specific services, thereby enhancing network performance and user experience.
1.1.2 More Intelligent Network Operation, Maintenance, and Management
@@
Compared to traditional core network equipment’s typical millisecond-level fault detection efficiency, network resource virtualization introduces new potential fault points. Moreover, it is difficult for operators to delineate the responsibility boundaries among suppliers when faults occur. Using software mechanisms for fault detection, the virtual network function at the edge site typically experiences longer response times, which complicates timely network maintenance and management. Furthermore, the continuous emergence of new applications and services demands higher bandwidth, lower latency, more excellent reliability, and enhanced security from the network. Traditional network operation and maintenance management methods will find it increasingly difficult to adapt to the future’s more complex and dynamic network environments. The integration of AI and wireless networks can support the in-depth analysis and mining of a large amount of network data, achieve accurate fault prediction and rapid diagnosis, significantly reduce the relatively long network fault recovery time, and enhance network availability.
@@ 2.1.2 Deep Learning
The principle of deep learning technology is to emulate the human brain [35], constructing artificial neural networks to learn knowledge from data. The neural layers can automatically learn and extract complex feature patterns from large-scale datasets, using gradient descent to update parameter errors. Table 4 summarizes some popular deep learning algorithms, and detailed descriptions are given as follows. Convolutional neural network (CCN) are widely used for image and video understanding [36]. Inspired by the functioning principles of the biological visual system, CNNs efficiently extract local features from image data via convolution operations and reduce computational complexity through pooling layers. Recurrent Neural Networks (RNN) represent a class of neural networks designed explicitly for sequential data within deep learning. Leveraging a recursive network structure[37], RNNs excel at capturing latent dependencies in sequential data. Generative adversarial networks (GANs) are a class of neural networks widely used for unsupervised learning. They comprise two neural networks, a discriminator and a generator, and use adversarial training to produce artificial data identical to actual data [38]. Deep Reinforcement Learning (DRL) is a convergence of deep learning and reinforcement learning[34], employing deep neural networks to approximate value functions or policy functions in reinforcement learning problems. This approach enables agents to autonomously learn and optimize decision-making processes to maximize long-term rewards in complex and dynamic environments[39, 40].
@@
CNN LeNet-5 Early convolutional neural network
@@ 3.2.2 AI-Based BS Energy Conservation
AI strategy can quickly analyze large-scale energy consumption data and discover hidden rules and correlations [84]. Therefore, network traffic can be accurately predicted according to dynamic data analysis results to reasonably plan parameters such as BS power and hibernation state and achieve accurate energy consumption control and resource optimization. The AI method can also systematically consider many factors, such as network coverage, capacity, and signal quality, to dynamically adjust the energy-saving strategy in the network and realize the global optimization decision. In addition, primary data such as the configuration and performance statistics of the live network can also be used to optimize the 5G device-level energy-saving function parameters based on feature clustering and neural network algorithms. Under the premise of ensuring service quality, the problem of 4G/5G overlapping coverage can be fully considered for the cell’s cooperative shutdown to maximize the effect of the device-level energy conservation function [85].
@@ 3.3.1 AI-Based Channel Estimation
Considering the sparse mmWave channel matrix as a natural image, the authors of [105] propose a practical and accurate channel estimation framework based on the fast and flexible denoising CNN. Simulation results validate that the training speed of a quick and flexible denoising CNN is more rapid than that of other channel estimators without sacrificing normalized mean square error performance. [106] considered the time-frequency response of a fast-fading communication channel as a 2D image. This scheme considers the pilot values as low-resolution images and uses an SR network cascaded with a denoising IR network to estimate the channel. The results confirm that this pipeline can be used efficiently in channel estimation. To address channel distortion, the model developed in [101] is first trained offline using the data generated from simulation based on channel statistics and then used for recovering the online transmitted data directly. [101] is more robust than conventional methods when fewer training pilots are used, the cyclic prefix is omitted, and nonlinear clipping noise exists. The authors of [107] exploited a learned denoising-based approximate message passing network, which can learn channel structure and estimate channels from many training data. Simulation results show that this neural network significantly outperforms state-of-the-art compressed sensing-based algorithms even when the receiver is equipped with a small number of RF chains.
@@ 3.3.2 AI-Based OFDM Receiver
The wireless communication transmitter and receiver are divided into multiple processing mods in OFDM. Each module is responsible for specific sub-tasks, such as Quadrature amplitude modulation, an IFFT, channel estimation, signal detection, etc. The advantage of this design is that each module can be analyzed and optimized individually, resulting in a very efficient and stable system.
@@
In [108], a model-driven deep learning (DL) approach was proposed that combines deep learning with expert knowledge to replace existing orthogonal frequency division multiplexing receivers in wireless communications. In [109], a machine learning (ML) assisted physical layer receiver technique was proposed to demodulate the OFDM signals, subject to very high Doppler effects and corresponding distortions in the received signal. To find a balance between full-size CP and non-existent CP, the authors of [110] investigate the redundancy problem and propose a minimum redundant OFDM receiver using deep learning (DL) tools. In [111], the receivers were designed based on deep neural networks (DNNS) in deep learning, consisting of a layer of DNNS and soft decisions. The problems of channel estimation error, delay, and decoding limits between users with traditional detection methods have been solved. It can recover the symbols of all users at one time and jointly perform channel estimation and signal detection. In [112], a novel generative supervised deep neural network is designed that can be trained using a reasonable number of pilots. After channel estimation, the neural network-based receiver (specifically, the autoencoder) jointly learns the pre-encoder and decoder for data symbol detection. In [113], an intelligent receiver for the OFDM communication system was designed based on the Densenet neural network structure, and the intelligent receiver is realized by optimizing the Densenet neural network structure. This method can recover information at the receiving end and avoid complicated pilot operations and signal error accumulation. In [114], a CNN (CNN) was used to reduce the bit error rate by using the mathematical calculation function of discrete Fourier transform (DFT) and the training of OFDM signal samples.
@@ 3.3.4 AI-Based CSI Feedback Algorithm
3.4 New Technological Applications Enabled
AI algorithms can dynamically allocate resources, such as bandwidth and power, to create and manage network slices in the wireless network. This can enable the creation of customized networks tailored to the needs of different applications and services. The booming development of AI has enabled the development of many new technological applications, such as network slicing, mobility management, and wireless localization.
@@ 4.2.1 Distributed Intelligence and Federated Learning
The NET4AI architecture must support low-latency, high-reliability communication computing requirements and increasing privacy concerns. Distributed intelligence and federated learning (FL) are the key technologies that can solve this problem. Traditional centralized architectures in wireless networks can no longer meet low-latency, high-reliability communication and computation demands. Additionally, conventional centralized architectures are inefficient in supporting the ubiquitous intelligence required by future wireless networks. This trend towards deploying intelligent decentralized elements has become apparent, with intelligence gradually shifting from centralized to distributed systems. Therefore, introducing distributed intelligent computing architectures is necessary to fully utilize the multi-dimensional data and computing resources held by user terminals and nodes, enabling intelligent connectivity.
@@
Figure 18: FL, Multi-Agent Reinforcement Learning and Split Learning
@@
Horizontal FL: Also known as sample-based FL, this can be applied in scenarios where the datasets of the various participants in FL have the same feature space but different sample spaces. Horizontal FL enhances privacy by training AI models on local devices and sharing only model parameters rather than raw data. Each node trains the model locally during training and periodically uploads model weights. The central node integrates these weights to form federated model weights, which are then sent back to the nodes for the next training cycle or local inference. FL ensures that user data remains local, transmitting only weights/gradients to achieve global training, thereby enhancing privacy.
@@
Through horizontal federation learning, different participants can jointly train a more powerful AI model based on local data sets while protecting privacy and improving the accuracy and generalization of the model. Additionally, horizontal FL enables real-time learning on distributed edge devices, allowing for timely model parameter updates to constantly changing environments and data. This is crucial for AI applications that require quick responsiveness and strong adaptability. It also supports collaboration across devices with varying computational capabilities and data distributions, catering to the diverse needs of 6G network terminals. Personalized model training on local devices enhances user experience by offering more individualized services.
@@
The main application areas of horizontal FL include smart healthcare, personalized services, and intelligent manufacturing[180]. The authors of [181] provide the framework for an advanced blockchain-orchestrated machine learning system for privacy-preserving FL in medicine and a new utility in health. The authors of [182] combined blockchain and FL to accelerate AI model training in healthcare and improve model robustness. The authors of [183] introduced personalized FL (PFL) methods with theoretical performance derivations, modeling user personalization and data distribution. The authors of [184] categorized PFL techniques based on key challenges and strategies, highlighting critical ideas and future research directions in PFL architecture, realistic benchmarks, and trustworthy methods. The authors of [185] propose a new FL framework for cross-domain prediction for intelligent manufacturing, designed to quickly adapt to new products, processes, and applications with limited training data.
@@
Vertical FL: Vertical FL refers to a setup where datasets have different features but share the same samples/users. It is suitable for scenarios where there is a high overlap of users among participants but low or no overlap in feature space by extracting parts where participants have the same users but not the same features for joint training. In the training process, it is first necessary to sample align each participant’s data to obtain overlapping user data, then train the model locally and upload the parameters to the server [186]. The difference between vertical and horizontal FL is that horizontal FL emphasizes cooperation and sharing between different data sources. In contrast, vertical FL focuses on collaboration and sharing within varying levels of the same data source.
@@
Vertical FL can leverage the correlation between different levels of data sources to achieve data augmentation. By sharing model parameters, various data sources can complement and enrich each other’s data, improving AI models’ generalization ability and robustness. Furthermore, vertical FL can facilitate cross-level learning and knowledge transfer. Different data sources can share model parameters and learning experiences, achieving cross-level knowledge transfer and migration and enhancing the overall model’s effectiveness.
@@
The main application areas of vertical FL include recommendation systems, smart healthcare, and smart transportation. The authors of [187] propose a method based on clustering and latent factor models under the vertical federated recommendation system. Considering the diversity of a large number of different users in each participant, users are clustered to reduce the dimension of the matrix and improve the accuracy of user recommendations. The authors of [188] describe the idea of Intention-Hiding Vertical FL and illustrate a framework to achieve this privacy-preserving goal. The authors of [189] propose a Federated Deep Learning based on the Spatial-Temporal Long and Short-Term Networks algorithm to predict traffic flow by utilizing observed historical traffic data. An additive homomorphic encryption method based on vertical FL trains local TFP models and shares short-term spatio-temporal hiding information.
@@
Split Learning: This technique separates machine learning model architecture between clients and servers, allowing joint model training without disclosing raw data. High-computation tasks are assigned to powerful central nodes, while shallower data processing layers are kept at data-holding terminals. Split learning offers better privacy than FL and reduces the processing load on resource-limited devices. Its applications focus on healthcare, where lightweight devices train intelligent models based on users’ private health data [196].